OpenStudy (anonymous):

Mindy conducted a survey of the phones in four stores and wrote the following observations: 5 of the 8 phones in Store 1 were smartphones 13 of the 22 phones in Store 2 were smartphones 22 of the 30 phones in Store 3 were smartphones 12 of the 25 phones in Store 4 were smartphones Based on the survey, which store had the greatest percentage of smart phones? Store 1 Store 2 Store 3 Store 4

OpenStudy (michele_laino):

Please you have to perform the least common denominator of the subsequent fractions: 5/8, 13/22, 22/30, 12/25

OpenStudy (anonymous):

I dont get it though! ;'(

OpenStudy (anonymous):

Which store is it!?!

OpenStudy (michele_laino):

otherwise, try to calculate each percentage, for esample 8/25=32%, 13/22=59%, 22/30=73,3% and 12/25 = 48%

OpenStudy (anonymous):

@pokamons

OpenStudy (anonymous):

@Hotchellerae21

OpenStudy (anonymous):

@kyrea

OpenStudy (anonymous):

i think it's store 3

OpenStudy (michele_laino):

@LIGHT_EM_UP In my preceeding answer I have made an error. I will go to remediate. for example: \[\frac{ 5 }{ 8 }=0.625 * 100=62.5 percentage\] \[\frac{ 13 }{ 22 }=0.59 *100 = 59 percentage\] \[\frac{ 22 }{ 30 }=0.733*100=73.3percentage\] \[\frac{ 12 }{ 25 }=0.48 *100 =48percentage\] now you have to find, please, which is the greaest percentage, please try

OpenStudy (griffindore):

hey

OpenStudy (griffindore):

sorry i just got back

OpenStudy (griffindore):

Sorry i dont no